1.00 mL of a 3.05 × 10–4 M solution of oleic acid is diluted with 9.00 mL of petroleum ether,
forming solution A. Then 2.00 mL of solution A is diluted with 8.00 mL of petroleum ether, forming solution B. How many grams of oleic acid are 5.00 mL of solution B? (Molar mass for oleic acid = 282 g/mol)
1
Expert's answer
2020-04-13T02:33:50-0400

Solution A

"Volume=10ml"

"n=\frac{3.05×10^{-4}}{10^{-3}}=3.05×10^{-7}" moles

Solution B

"Volume=10ml"

"n=3.05×10^{-7}×\frac{2}{10}=0.61×10^{-7}"

Moles of oleic acid in "5ml" of Solution "B=0.61×10^{-7}×\frac{5}{10}=0.305×10^{-7}moles"

Mass of oleic acid is "0.305\u00d710^{-7}\u00d7282=8.601\u00d710^{-6} g"